Oxygen gas having volume of 500cm3 at 27 degree and pressure of 8.31Pa .expands until its volume becomes 1.0m3 and its pressure 10Pa.Find the mass in moles of oxygen and its final temperature.
1
Expert's answer
2020-03-03T10:03:58-0500
  • Settle through the Combined Gas Law:
"{p_1V_1 \\over T_1}\t= {p_2V_2\\over T_2}"

"T_2 = {T_1p_2V_2 \over p_1V_1}"

"T_2 = {27*10*1 \\over 8.31*0.5}\t= 64.98 degree"

  • According to the Law of Mendeleev-Clapeyron:
"n = {pV \over RT}"

We assume that the temperature in the condition of the problem is indicated in degrees Kelvin:



"n = {10*1 \\over 8.314*64.98}\t = 0.0185mol"

"m = n*M = 0.0185*32 = 0.59g"